OpenStudy (anonymous):

What is the y-intercept of the line? Y=-2x-8 y-intercept = Is it -2?

OpenStudy (anonymous):

Nah it's actually -8. Because the y-intercept is what y will equal when x = 0. y = -2x - 8 y = -2(0) - 8 y = -8. Or you could just look at the number that isn't being multiplied by x. -8.

OpenStudy (anonymous):

ok